OpenStudy (anonymous):

Medal PLease Write an equation of a line parallel to line GH below in slope-intercept form that passes through the point (-5, 6). Line GH is shown. G is at negative 2, 2. H is at 2, 6. y = −x − 1 y = x − 1 y = −x + 11 y = x + 11

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (anonymous):

@TwinYang

OpenStudy (anonymous):

twin can you give me the answer then the explanation reason is is bcuz i only have 30 mins

OpenStudy (anonymous):

please

OpenStudy (anonymous):

ill medal you too

OpenStudy (anonymous):

ninja please my grade depends on it

OpenStudy (anonymous):

i have like 5 questions

OpenStudy (anonymous):

Formula is y= mx +b m = slope of the line b = the y-intercept

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

i have no time to work these out thats what im trying to say

OpenStudy (anonymous):

i really need some quick help pronto

OpenStudy (anonymous):

anyone please i dont want to fail this is really important

OpenStudy (anonymous):

First we need to find the slope of line GH. Here's the line graphed. What's the equation of this line?

OpenStudy (anonymous):

idk how to find it

OpenStudy (anonymous):

What's the y intercept?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ninja i dont have time for this man

OpenStudy (anonymous):

What's the slope?

OpenStudy (anonymous):

im gonna fail bro

OpenStudy (anonymous):

slope is 4

OpenStudy (anonymous):

i dont want to fail man

OpenStudy (anonymous):

please just answers i will medal whoever can come up with them explain later please

OpenStudy (anonymous):

this is really important

OpenStudy (anonymous):

y = x + 11 Use this to help you find the slope-intercept http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html

OpenStudy (anonymous):

Yup...that's the answer...

OpenStudy (anonymous):

y = x + 11 passes through the point (-5, 6).

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

For the line segment whose endpoints are L (0, 1) and M (2, 8), find the y value for the point located 3 over 5 the distance from L to M. 5.2 3.5 4.8 1.6

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (anonymous):

@TwinYang

OpenStudy (anonymous):

do you know it?

OpenStudy (anonymous):

nope i suggest you make it a new question so people can help.

OpenStudy (anonymous):

ok